OpenStudy (anonymous):

Can someone please help me with this 10 question quiz? I suck at math. ):

OpenStudy (anonymous):

okaiii hit us

OpenStudy (anonymous):

The relation {(-5, 5), (-6, 7), (-7, 7), (-8, 0)} is a function. TRUE FALSE

OpenStudy (anonymous):

what do you think it is???

OpenStudy (anonymous):

look at your x values, are there 2 of the same number? that have 2 outputs?

OpenStudy (anonymous):

No... So does that mean that it's true..?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

look to see if the first numbers repeat if they do, NOT a function if they are all different, then it IS a function ignore the second numbers

OpenStudy (anonymous):

okay next one

OpenStudy (anonymous):

What is the range of the inverse of relation {(1, 7), (2, -4), (5, 6), (2, 8)}? {1, 2, 5} {-4, 6, 7, 8} {1, 5} {-4, 7, 8}

OpenStudy (anonymous):

Oh and thank you so much! (:

OpenStudy (anonymous):

do you know what domain and range is?

OpenStudy (anonymous):

Domain is x range is y

OpenStudy (anonymous):

so your answer would be?

OpenStudy (anonymous):

Oh... so then it's B?

OpenStudy (anonymous):

Example 1 is the domain 7 is the range get it?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

Find the inverse of f(x) = 2x - 2. f-1(x) = -2x + 1 f-1(x) = -2x + 2 [DAM:NONE]

OpenStudy (anonymous):

1 sec just double checking my work

OpenStudy (anonymous):

I know but I dropped out when I was a freshman and now I'm 17 and trying to get my diploma. :P I'm just not that great at school. I barely past through middle school.

OpenStudy (anonymous):

your first step is to complete the square which should give you y=x^2-2x

OpenStudy (anonymous):

How did you do that?

OpenStudy (anonymous):

now bring your 1 to both sides y+1=x^2-2x+1 and just sqrt the 2 so it will have your leftover 1

OpenStudy (anonymous):

and by the looks of my work i dont think this is a function

OpenStudy (anonymous):

not a function

OpenStudy (anonymous):

Okay so then it would be C?

OpenStudy (anonymous):

correctt

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) - g(x). 5x + 4 x - 4 x + 14 5x - 4

OpenStudy (anonymous):

okay so this one is pretty much plugging in the answers

OpenStudy (anonymous):

they tell you that f(x)=3x+5 and that g(x)=2x-9

OpenStudy (anonymous):

So how would you do that?

OpenStudy (anonymous):

they want you to plug those in to this problem----->f(x) - g(x)

OpenStudy (anonymous):

Oh okay.. So let me work it out to see if I get it really quick...

OpenStudy (anonymous):

so it would be----->3x+5-2x-9

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Okay I got it, so now what would you do after that?

OpenStudy (anonymous):

combine your like terms

OpenStudy (anonymous):

show me what you get

OpenStudy (anonymous):

I got x-4

OpenStudy (anonymous):

there you go! lol doing good :P

OpenStudy (anonymous):

next oneee

OpenStudy (anonymous):

If f(x) = x - 4 and g(x) = 2x - 3, find g and f

OpenStudy (anonymous):

@Jelia97 after you get done with helping her, can you help me please?

OpenStudy (anonymous):

2x + 3 2x + 7 2x - 7 2x - 11

OpenStudy (anonymous):

so what they want you to do is put the 2 equations together

OpenStudy (anonymous):

yes i will ari

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

all you gotta do is combine them, what do you get?

OpenStudy (anonymous):

Okay so do I include the f(x) and the g(x)

OpenStudy (anonymous):

nope just put them together (x - 4 )( 2x - 3) combine like terms

OpenStudy (anonymous):

oh okay. I got 2x-7

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

do you think you can handle the rest of your quiz?

OpenStudy (anonymous):

Yes. Thank you so much for your help! I really appreciate it! <3 (:

OpenStudy (anonymous):

welcome! good job :D

OpenStudy (anonymous):

ill still be on if you need help tho lol

OpenStudy (anonymous):

lol Thanks.